OpenStudy (braydonlevi99):

hours, and the second mechanic worked for 5 hours. Together they charged a total of $1075 . What was the rate charged per hour by each mechanic if the sum of the two rates was

OpenStudy (braydonlevi99):

My bad here's the question and I promise it makes sense: Two mechanics worked on a car. The first mechanic worked for 10 hours, and the second mechanic worked for 5 hours. Together they charged a total of $1075. What was the rate charged per hour by each mechanic if the sum of the two rates was $140 per hour?

jhonyy9 (jhonyy9):

first mechanic let note FM second mechanic note SM FM worked 10 h SM ----- 5 h ----------------- together 15 h in what they charged 1075 $ - because the sum of the two rates of charge was 140 $/h and because they worked together 5 hours so 5*140=700$ - what mean that this 700$ divided by 2 result have got 350$-350$ for 5 hours the work - so divided 350 by 5 = 70$ per hours - and like a proof 70+70=140$/h what they have charged together - than result that the difference 1075-700 = 375$ was charged just the FM alone - hope this will help you -